mathematics written work are normally distributed with a mean of 23 and standard deviation of 9. describe the sampling distribution of 15 students who are normally from this population



1
Expert's answer
2022-04-22T12:59:45-0400

"\mu_{\bar X}=\mu=23."

"\sigma_{\bar X}=\frac{\sigma}{\sqrt{n}}=\frac{9}{\sqrt{15}}=2.32."